OpenStudy (anonymous):

Find the point, M, that divides segment AB into a ratio of 2:1 if A is at (-1, 2) and B is at (8, 15).

OpenStudy (anonymous):

is it A) y= -bx-b(1-a) B) y=bx-b(a-1)

OpenStudy (anonymous):

C) \[y=1/bx+b(a-1)\]

OpenStudy (anonymous):

or D) Y=-1/bx+b(1-a)

OpenStudy (anonymous):

wrong